# Dynamical Phase Transitions, Caustics, and Quantum Dark Bands

### v1

@@ VI Switching Lines in Fock Space
In the FCTFIM the DPT is merely a single point of a more general switching line where nonanalytic behavior occurs due to overlapping exponential tails. In our example Fig. 2 we expect a switching line to originate at the intersection of the two caustics (that separate the $k=0$ and $k=1$ trajectory classes), and passing through the DPT point. We confirm this expected behavior numerically and the resulting switching line is shown in Fig. 9.
@@
VII Discussion and Conclusions
By explicitly calculating the Loschmidt echo in the semiclassical regime using the Poisson summation formula, we have shown that DPTs (or more precisely, type-II DPTs in the classification scheme given in reference [25]) in the FCTFIM arise due to the switching in exponential dominance between evanescent tails originating from two neighbouring fold caustics in Fock space. This switching occurs as a function of time and is essentially the quantum analogue of the switching that occurs as a function of angle between the primary and secondary bows on either side of Alexander’s DB in ordinary rainbows seen in the sky. We therefore have a striking analogy between the behavior of a microscopic quantum system and a macroscopic ‘everyday’ natural phenomenon. Motivated by this analogy, we dub the period of time during which the Loschmidt echo is made up of only exponentially small tails as a quantum DB.
@@ VII Discussion and Conclusions
Since caustics are a common feature of nonequilibrium classical and quantum waves it would seem likely that the quantum DB mechanism proposed in this paper has wider applicability than the FCTFIM. However, the FCTFIM is an integrable model, as is the 1D TFIM which was the original setting of DPTs [26]. The question then arises as to whether quantum DBs can also occur in nonintegrable models? Indeed, caustics arise from the projection down onto subspaces such as the momentum or coordinate planes of the motion around the invariant tori that foliate phase space in classical integrable systems [103]. The invariant tori themselves occur due to the complete set of conserved quantities that exist in integrable systems, and are therefore destroyed by strong chaos where these conserved quantities are absent. However, between complete integrability and strong chaos there lies the much more typical regime of weak chaos. The celebrated Kolmogorov-Arnold-Moser (KAM) theory shows that some invariant tori survive in this intermediate regime [104], and hence caustics do too. In a recent paper on caustics in Heisenberg spin chains it was shown that caustics do indeed have structural stability against weak chaos which was introduced by adding non-integrable terms to the Hamiltonian in a controlled way [102], and this result is in accordance with the possibility of a quantum version of the KAM theory that would preserve wave catastrophes in weakly nonintegrable regimes [105]. Moreover, DPTs have been specifically shown to also occur in nonintegrable models [106]. Thus, although further investigation is needed, there is substantial evidence suggesting that the quantum DB mechanism can also be used to explain DPTs in weakly chaotic systems. A particularly interesting case for future investigation would be whether quantum DBs can be seen in many-body localized states because these are characterized by a full set of locally conserved quantities and so have some similarties to integrable systems, but at the same time are disordered [107].
@@
The interpretation of the DPT in terms of evanescent tails coming from different caustics allowed us to propose a measurement scheme that could be used to experimentally verify the existence of a DPT in this system. The idea is to restrict the quantum state to a local region in phase space that entails a single caustic when projected onto Fock-space. This way we can isolate the amplitude contributions due to the specific caustic. In a real system one can realize this by considering a bipartition of the system and conditioning a Fock-state measurement of the first subsystem onto a phase-space position measurement on the second subsystem. This makes it possible to measure separately the Airy-like functions that dress the caustics and cross in the quantum DB region. Such a nontrivial protocol could be realized with state-of-the-art quantum simulators, for instance trapped ion platforms [30]. An open question in this regard is whether an observation of singular features in the many-body wave function such as DPTs could also be made visible in large collective systems, which is the natural setting for cavity-QED with cold atoms [55].
@@ Acknowledgements.
DO acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) [Ref. No. RGPIN-2017-06605].
@@ Appendix B Analytical continuation via complex mechanics
Note that by fixing the initial and final conditions one recovers the WKB action $S(E,\eta)$ from appendix IV, where the energy $E$ has to be chosen in correspondence to the time $t$. In the classically forbidden region no classical trajectories exist that connect the given initial condition with the given final condition at time $t$, and we expect the wave function to be exponentially decaying. Consequently, $S(t,\eta)$ will no longer be real-valued. In order to still solve the Hamilton-Jacobi equation one can consider complex initial conditions $\phi_{cpx}(0)\in\mathbb{C}$ and search for complex solutions of Hamilton’s equations satisfying the boundary condition $\eta_{cpx}(0)=\eta_{0}\in\mathbb{R}$ and $\eta_{cpx}(t)=\eta\in\mathbb{R}$. This leads to both complex energies and a complex action that can be obtained from equation (54), upon replacing $\phi_{cl},\,\eta_{cl}$ with $\phi_{cpx},\,\eta_{cpx}$. We indeed find that at any time in the forbidden region two complex trajectories exist with a positive imaginary part of the action corresponding to an exponentially small wavefunction, see Fig. 10. These two families of trajectories originate from the first and second caustic respectively, i.e. they are continuously connected to the classical trajectories on the bright side of the corresponding caustic. Since we work directly in a time-dependent framework, the action (54) obeys the steepest descent condition (25) by construction. Thus, it is not necessary to compute energy-dependent actions for arbitrary complex energies $E$. The imaginary parts of the time-dependent complex actions directly yield the exact asymptotic rate function as displayed in Fig. 4 (right panel, blue and red curves).

### v2

@@ VIII Switching Lines in Fock Space
In the FCTFIM the DPT is merely a single point of a more general switching line where nonanalytic behaviour occurs due to overlapping exponential tails. In our example Fig. 2 we expect a switching line to originate at the intersection of the two caustics (that separate the $k=0$ and $k=1$ trajectory classes), and passing through the DPT point. We confirm this expected behaviour numerically and the resulting switching line is shown in Fig. 9.
@@
[…]
IX Discussion and Conclusions
By explicitly calculating the Loschmidt echo in the semiclassical regime using the Poisson summation formula, we have established a novel interpretation of the origin of DPTs (or more precisely, type-II DPTs in the classification scheme given in reference [25]) in the FCTFIM via the switching in exponential dominance between evanescent tails originating from two neighbouring fold caustics in Fock space. This switching occurs as a function of time and is essentially the quantum analogue of the switching that occurs as a function of angle between the primary and secondary bows on either side of Alexander’s DB in ordinary rainbows seen in the sky. We therefore have a striking analogy between the behaviour of a microscopic quantum system and a macroscopic ‘everyday’ natural phenomenon. Motivated by this analogy, we refer to the period of time during which the Loschmidt echo is made up of only exponentially small tails as a quantum DB.
@@ IX Discussion and Conclusions
The interpretation of the DPT in terms of evanescent tails coming from different caustics allowed us to propose a measurement scheme that could be used to experimentally verify the existence of a DPT in this system. The idea is to restrict the quantum state to a local region in phase space that entails a single caustic when projected onto Fock-space. This way we can isolate the amplitude contributions due to a specific caustic. In a real system one can realize this by considering a bipartition of the system and conditioning a Fock-state measurement of the first subsystem onto a phase-space position measurement on the second subsystem. This makes it possible to measure separately the Airy-like functions that dress the caustics and cross in the quantum DB region. Such a nontrivial protocol could be realized with state-of-the-art quantum simulators, for instance trapped ion platforms [38]. An open question in this regard is whether an observation of singular features in the many-body wave function such as DPTs could also be made visible in large collective systems, which is the natural setting for cavity-QED with cold atoms [61].
Since caustics are a common feature of nonequilibrium classical and quantum waves it would seem likely that the quantum DB mechanism proposed in this paper has wider applicability than the FCTFIM. However, the FCTFIM is an integrable model, as is the 1D TFIM which was the original setting of DPTs [26]. The question then arises as to whether quantum DBs can also occur in nonintegrable models? Indeed, caustics arise from the projection down onto subspaces such as the momentum or coordinate planes of the motion around the invariant tori that foliate phase space in classical integrable systems [107]. The invariant tori themselves occur due to the complete set of conserved quantities that exist in integrable systems, and are therefore destroyed by strong chaos where these conserved quantities are absent. However, between complete integrability and strong chaos there lies the much more typical regime of weak chaos. The celebrated Kolmogorov-Arnold-Moser (KAM) theory shows that some invariant tori survive in this intermediate regime [108], and hence caustics do too. In a recent paper on caustics in Heisenberg spin chains it was shown that caustics do indeed have structural stability against weak chaos which was introduced by adding non-integrable terms to the Hamiltonian in a controlled way [106], and this result is in accordance with the possibility of a quantum version of the KAM theory that would preserve wave catastrophes in weakly nonintegrable regimes [109]. Moreover, DPTs have been specifically shown to also occur in nonintegrable models [110]. Thus, although further investigation is needed, there is substantial evidence suggesting that the quantum DB mechanism can also be used to explain DPTs in weakly chaotic systems.
It might also be wondered whether quantum caustics can occur in mixed states and hence survive the effects of decoherence inherent in an open quantum system? The answer is yes: the fate of caustics has been investigated in a system equivalent to the FCTFIM studied here but subject to a continuous measurement of the magnetization such that the measurement-induced backaction leads to decoherence [75]. Solutions of the master equation for the density matrix reveal that caustics are still present at short times although at long times they eventually get washed away by the noise inherent to a quantum measurement. As might be expected, the rate of attrition depends on the strength of the measurement and thus the caustics last longer when the measurement is weak (the backaction in fact avoids a paradox whereby measurements could introduce singularities via the caustics [75]).
There are potentially many other systems and situations in which the quantum caustic mechanism for DPTs proposed in this paper might be tested in future work, but two seem particularly pertinent. The first type are spatially extended systems with short range interactions, such as the ordinary TFIM on a spin chain, where Hilbert space is larger than the FCTFIM and a desciption in terms of a collective macroscopic (classical) state is less applicable. For this case it is worth pointing out that caustics are known to occur in real space in the 1D TFIM and in fact they provide an explanation for the light cone-like transmission of information through the system following a quench [66]. A similar study of quenches in 1D Heisenberg spin chains revealed that caustics also occur in out-of-time-ordered correlators (OTOCs), underlining the generic nature of the phenomenon. We therefore speculate that caustics, being generic dynamical singularities in both the wave function and correlators, will continue to play a key role in the dynamics of systems with larger Hilbert spaces. Catastrophe theory guarantees the existence of higher dimensional caustics (see Ref. [67] for three dimensional caustics in quantum many-body dynamics), although the calculations in this case will be harder due to the higher complexity and it may be more tractable to switch to a statistical version of catastrophe theory [46].
A second interesting future direction concerns whether quantum DBs can be seen in many-body localized states in disordered systems. These are characterized by a full set of locally conserved quantities and so have some similarties to integrable systems where constants of the motion are responsible for caustics, and yet are different because the conserved quantities in the latter are global not local [111].
@@ Acknowledgements.
We are grateful to Professors M. V. Berry, D. M. Gangardt and C. J. Howls for discussions concerning the Stokes phenomenon, and Professor Markus Oberthaler for his question during a seminar as to whether there might be a connection between quantum catastrophes and DPTs, which provided one of the motivations behind this work. DO acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) [Ref. No. RGPIN-2017-06605].
@@ Appendix B Analytical continuation via complex mechanics
Note that by fixing the initial and final conditions one recovers the WKB action $S(E,\eta)$ from appendix VI, where the energy $E$ has to be chosen in correspondence to the time $t$. In the classically forbidden region no classical trajectories exist that connect the given initial condition with the given final condition at time $t$, and we expect the wave function to be exponentially decaying. Consequently, $S(t,\eta)$ will no longer be real-valued. In order to still solve the Hamilton-Jacobi equation one can consider complex initial conditions $\phi_{cpx}(0)\in\mathbb{C}$ and search for complex solutions of Hamilton’s equations satisfying the boundary condition $\eta_{cpx}(0)=\eta_{0}\in\mathbb{R}$ and $\eta_{cpx}(t)=\eta\in\mathbb{R}$. This leads to both complex energies and a complex action that can be obtained from equation (55), upon replacing $\phi_{cl},\,\eta_{cl}$ with $\phi_{cpx},\,\eta_{cpx}$. We indeed find that at any time in the forbidden region two complex trajectories exist with a positive imaginary part of the action corresponding to an exponentially small wave function, see Fig. 10. These two families of trajectories originate from the first and second caustic respectively, i.e. they are continuously connected to the classical trajectories on the bright side of the corresponding caustic. Since we work directly in a time-dependent framework, the action (55) obeys the steepest descent condition (26) by construction. Thus, it is not necessary to compute energy-dependent actions for arbitrary complex energies $E$. The imaginary parts of the time-dependent complex actions directly yield the exact asymptotic rate function as displayed in Fig. 4 (right panel, blue and red curves).